OpenStudy (anonymous):

Can you help me with this derivative? f(x)=ln(3x)

OpenStudy (anonymous):

using u-substitution, u = 3x du = 3dx f(x) = ln(u) f'(x) = u'/u f'(x) = 3dx/3x f'(x) = dx/x

OpenStudy (anonymous):

You should be in a Math forum but here's the answer anyway. If you're using the u substitution, you should remember the chain rule, so u = 3x u' = 3 f'(x) = d/du[ln(u)] * u' f'(x) = (1/3x)*3 f'(x) = 1/x Another solution: f(x) = ln(3x) By applying logarithm rules, we get f(x) = ln(3) + ln(x) f'(x) = d/dx [ln(3)] + d/dx [ln(x)] f'(x) = 0 + 1/x f'(x) = 1/x